A survey of MAT 101 and 102 students showed 94 owned laptops, 19 owned a laptop and a calculator and 9 owned only a calculator. With the entire number of students surveyed being 185, find the probabilities that a randomly selected MAT 101/102 student:
a)Does not own a laptop or a calculator.
b)Does not own a laptop.
c)Owns a laptop or a calculator, but not both.

Answer by stanbon(75887)

A survey of MAT 101 and 102 students showed 94 owned laptops, 19 owned a laptop and a calculator and 9 owned only a calculator. With the entire number of students surveyed being 185, find the probabilities that a randomly selected MAT 101/102 student:
====
Sketch a Venn Diagram with 2 intersecting circles inside a rectangle.
label one circle "laptop" with 94
write 19 in the intersection
label the interior of the other circle "9" and total in that circle = 9+19 = 28
label the rectangle with "185"
-------
a)Does not own a laptop or a calculator.
Ans:: (185-(94+9))/185 = 82/185
-------------------------
b)Does not own a laptop.
Ans:: (185-(9+82))/185 = 94/185
-------------------------
c)Owns a laptop or a calculator, but not both.
Ans: (75+9)/185 = 84/185
------------
Cheers,
Stan H.